OpenStudy (anonymous):

Indicate which property is illustrated in Step 3. Step 1 8x + 2 + 6x + 5 = (8x + 2) + (6x + 5) Step 2 = 8x + (2 + 6x) + 5 Step 3 = 8x + (6x + 2) + 5 Step 4 = (8x + 6x) + (2 + 5) Step 5 = (8 + 6)x + (2 + 5) Step 6 = 14x + 7 A. arithmetic fact B. commutative C. distributive D. associative

OpenStudy (bee_see):

b

OpenStudy (bee_see):

it means to swap like a+b=b+a, so 2+6x = 6x+2

OpenStudy (triciaal):

agree